Assume that no denominator equals 0:
10^8/10^5

Answer by checkley71(8403)

10^8/10^5=10^3 ANSWER.
PROOF
10*10*10*10*10*10*10*10/10*10*10*10*10 NOW CANCEL OUT 5 OF THE 10s IN BOTH THE NUMERATOR & THE DENOMINATOR.
10*10*10=10^3